Use the definition of limit to prove that the sequence {n − 1 }∞n=1 is divergent.


1
Expert's answer
2021-01-26T19:05:11-0500

"Solution: Using ~the ~definition~ we~have, Let~a_n={n-1}
\\ \lim_{n \to \infty} a_n = \lim_{n \to \infty} {n-1}=L
\\there~ exists~ an ~\epsilon~such~that~for~all~N>0,there~is~an~n>N~with~
\\|a_n-L| \geq ~\epsilon~~~~~................................................(1)
\\Since~ when~ we~ put~n=1,2,3,4,.......
\\we~get~sequence~\{-1,0,1,2,3,4,......\} 
\\ when ~we~put~\lim_{n \to \infty} ~it ~approches~ to ~\infty. 
\\\therefore , by ~(1),~sequencce~\{a_n\} ~diverges."